Question 2 The following is a list of history midterm scores. Incorrect 74 83 70 76 0.00 points out of 5.00 75 59 92 56 50 Flag 57 70 52 74 question 83 78 13 51 70 57 Find the five number summary for the data. (Put them in as min,Q1,Median,Q3,max with no spaces. For example; 3,6,7.5,9,14) Answer 13,55,70,77,92 Question 3 Find the mean and standard deviation for the data of the previous problem. Put Partially correct them in the order mean,standard deviation (no spaces) rounded to the nearest tenth. For example 15.2,19.7 2.50 points out
Question The frequency table is shown for the age of fire fighters in the city of Cleansburg Correct 27 37 Age 25 28 29 30 31 32 5.00 points out Freq. 2 15 12 of 5.00 Flag a)What percentage of firefighters are 33 and older? (Round to the nearest percent.) question b)How many degrees of a pie chart would this take? (Round to the nearest degree.) Input answers as percent,degrees again no spaces (don't put % or degree symbol, just the numbers.) Answer 20,72 Question 5 Find the mean for the data of the previous problem. (Round to the nearest tenth.) Incorrect 0.00 points out Answer7.5 of 5.00
Answer #1

question 2)

Sr. no sorted X
1 13
2 50
3 51
4 52
5 54
6 55
7 55
8 55
9 56
10 57
11 57
12 59
13 70
14 70
15 70
16 74
17 74
18 75
19 76
20 77
21 77
22 78
23 83
24 83
25 92

Median=0.5(n+1)th value =    13   th value of sorted data
=   70  
      
quartile , Q1 = 0.25(n+1)th value=   6.5   th value of sorted data
=   55  
      
Quartile , Q3 = 0.75(n+1)th value=   19.5   th value of sorted data
=   76.5  
      
IQR = Q3-Q1 =    21.5  
      
1.5IQR =    32.25  
      
lower bound=Q1-1.5IQR=   22.75  
      
upper bound=Q3+1.5IQR=   108.75  
      
outlier =values outside lower bound and upper bound      
count below lower bound=   1  
count above upper bound=   0  
total outlier =    1  

maximum =    92

five number summary is (50,55,70,76.5,92)

-----------------

question 3)

Sr. no X (X - X̄)²
1 74 89.870
2 75 109.830
3 57 56.550
4 83 341.510
5 77 155.750
6 83 341.510
7 59 30.470
8 70 30.030
9 55 90.630
10 55 90.630
11 70 30.030
12 92 755.150
13 52 156.750
14 78 181.710
15 51 182.790
16 55 90.630
17 56 72.590
18 74 89.870
19 13 2654.310
20 70 30.030
21 76 131.790
22 50 210.830
23 77 155.750
24 54 110.670
25 57 56.550
X (X - X̄)²
total sum 1613 6246.240
n 25 25

mean =    ΣX/n =    1613.000   /   25.000   =   64.5
                      
sample variance =    Σ(X - X̄)²/(n-1)=   6246.2400   /   24.0000   =   260.2600
                      
sample std dev =   √ [ Σ(X - X̄)²/(n-1)] =   √   260.2600   =       16.1

answer: 64.5 , 16.1

---------------------------------------------------------------------------------------------------